OpenStudy (anonymous):

A D E G J three letters are to be picked form the letters listed above. if the second letter must be a vowel and the first or third letter must be G, how many different if no letters can be repeated? a. 12 b. 14 c. 16 d. 18 e. 20

OpenStudy (anonymous):

are you sure you paste the right multiple choices for this question? I only find 6

OpenStudy (anonymous):

omg, never mind. I though A was a consonant. answer is 12. 1 * 2 * 3 = 6 3 * 2 * 1 = 6 6 + 6 = 12